# SAM 2++: Tracking Anything at Any Granularity

### v2

@@ 1 Introduction
Video tracking has been a fundamental task in computer vision for decades, aiming to estimate the state of an arbitrary target in video sequences given its initial status. Despite sharing this core objective, the tracking domain has fragmented into several independent sub-tasks based on different target granularities, including Single Object Tracking [23, 44, 31] (SOT) with bounding box, Video Object Segmentation [48, 60, 28] (VOS) with precise pixel-level mask, and Point Tracking [5, 71, 21] with tiny points. This fragmentation based on state granularity has led most video tracking research to focus on a specific task and propose specialized designs only for that task. While this design trend enhances tracking performance, it limits the generalization ability of tracking models across multiple tasks and results in redundancy in both model design and parameters. To unify tasks, current unified vision models typically share feature extraction backbones while employing task-specific branches [73], convert those tasks into a seq2seq framework [8], or share one appearance model for either propagation or association [55, 69, 63, 64, 53]. However, they choose to provide different interfaces for different tasks, rather than seeking a unified visual representation of tracking targets, and ignore the point tracking task.
Unlike them, we observe that these seemingly disparate tracking paradigms fundamentally differ primarily in their state granularity, while sharing the memory matching strategy: the model encodes the previous state into memory, and matches the current features with the stored memory when a new frame is received. Based on this strategy, we decide to unify target states at three different granularities through a uniform memory representation. Recently, Segment Anything Model 2 [50], a strong foundational model, has been proposed for high-quality video object segmentation given various prompts. Due to its flexible prompt mechanism and powerful mask tracking capabilities, we extend this model to track arbitrary granularity, termed as SAM 2++.
@@ 2 Related work
Segment Anything Model. SAM [37] is a foundational model for high-quality segmentation given various prompts, and SAM 2 [50] extends it to video with streaming memory, effectively handling motion and occlusion, and they inspire many variant models In the image domain, HQ-SAM [35] enhances segmentation quality through a High-Quality Token, SAMRefiner [42] improves fine-grained details via a noise-tolerant prompt, CAT-SAM [59] adopts a conditional tuning approach to adapt to specialized image domains, and SAM-Adapter [9] incorporates lightweight adapters for improved downstream performance. In the video domain, SAM2Long [20] employs constrained tree search to reduce error accumulation, SAMURAI [65] uses the Kalman filter to select motion-aware memory, while DAM4SAM [52] introduces a distractor-aware memory. SAMWISE [17] and AL-Ref-SAM-2 [32] add additional prompts for more referring tasks. Despite these advances, these works remain task-specific, lacking cross-domain generalization and requiring separate implementations for each application.
Unified Vision Models. Recent years have witnessed significant progress in developing unified vision models that handle multiple tasks through shared architectures and demonstrate strong generalizability and flexibility. Pix2Seq [8] reformulates vision tasks as sequence generation problems, Uni-Perceiver [73] establishes unified representation spaces across modalities with shared encoders and decoders. UniTrack [55] demonstrates that video tracking tasks can be solved by a single appearance model with task-specific heads, while Unicorn [63] and UNINEXT [64] unify various tracking paradigms through common frameworks with different representations. Despite their impressive capabilities, these unified approaches predominantly focus on object-level tasks while neglecting finer-grained tasks such as point tracking. Furthermore, they do not take into account unifying video tracking tasks with various granularities through a unified visual representation.
@@ 3 Preliminaries: Segment Anything Model 2
The Segment Anything Model (SAM) [37] is a milestone vision foundation model for class-agnostic image segmentation. It flexibly handles various prompts (box, point, mask) by encoding them into a unified embedding and has established an iterative data engine with model-assisted labeling to address dataset limitations. SAM 2 [50] extends SAM to promptable video segmentation by introducing a streaming memory that stores previous target information and predictions. It comprises four main components: (i) a hierarchical image encoder that encodes each frame ${I}_{img}$ into image embeddings $F_{img}$, (ii) a prompt encoder, (iii) a memory mechanism (memory encoder, memory bank, memory attention), and (iv) a mask decoder for prediction.
@@
Mask Decoder. The mask decoder takes prompt embedding $\mathcal{P}_{sparse}$ and $\mathcal{P}_{dense}$, memory-conditioned image embeddings $\bar{F}_{img}\in\mathbb{R}^{C/4\times H/16\times W/16}$ (which we will explain latter), and a set of learnable tokens $\mathcal{E}_{tokens}$ as inputs. The learnable tokens contain an existence token $\varepsilon_{obj}\in\mathbb{R}^{C}$ to predict whether the target exists, an IoU token $\varepsilon_{iou}\in\mathbb{R}^{C}$ to predict the result accuracy, and multiple mask tokens $\varepsilon_{mask}^{N}\in\mathbb{R}^{N\times C}$ used to obtain $N$ mask candidates. To fuse the prompt embedding, a Two-Way Transformer $\mathbf{twTrans}$ [50] processes them as:
@@ 4.1 Unified Task Input and Output Processing
Input Unification via Task-Specific Prompt. Due to the input of the three tracking tasks having inconsistent granularity, we first unify inputs with task-specific prompts for different tasks. The video object segmentation task still adopts mask input ${I}_{box}$ as its prompt in mask form, and the single object tracking task takes its box input ${I}_{box}$ as the prompt. As for the point tracking task, expect the point coordinates ${I}_{point}$, we add a dense mask ${G}_{point}$ for additional prompt as a Gaussian map centred on the point and parameterised by sigma $\sigma$ and radius $r$ as: $G_{point}=\exp\left(-\frac{\|p-p_{0}\|^{2}}{2\sigma^{2}}\right)\cdot\mathbf{1}_{\{\|p-p_{0}\|\leq r\}}$ to highlights the point in mask form, maintaining consistency with output from Unified Decoder and source for Memory Encoder, which is better than naive {0, 1} mask. More importantly, we gradually decrease the radius and sigma during training to facilitate smoother convergence and more stable learning.
Output Unification via Unified Decoder. To unify the output of various tasks, we extended the Mask Decoder of SAM 2 as Unified Decoder, which also processes memory-conditioned image embeddings, prompt embeddings, and learnable tokens. For the SOT task, the outer box of the mask output $M^{N}_{box}$ cannot be used as task output because the complexity of the mask reduces the accuracy of the box, which focuses on the center point’s position and target scale. Instead, we add a Corner-based Head [62], $\mathbf{CornHead}$, which explicitly optimizes for the accuracy and stability of the bounding box and is widely used in the SOT task, to produce box predictions. As for the PT task, rather than direct point coordinates, we obtained point predictions in terms of mask predictions by soft-argmax operation during training or argmax operation during inference. This specific design aims for the output of the point task to be consistent with the source of memory, thereby achieving a unified encoding mask output for memory information of different granularities, and also helps to optimize model training. In summary, our unified decoder can be written as:
@@ 4.2 Task-adaptive Memory
Therefore, we propose a task-adaptive memory mechanism, which relaxes the uniformity by decoupling only the memory components: each task has its own convolutional Memory Encoder, and each applies an independent LoRA [30] in the transformer-based Memory Attention. This decoupled design effectively meets the diverse needs of different tasks and avoids the performance drop seen in a fully parameter-shared model, while keeping the overall structure consistent, enhancing the multi-task processing capability of the model. Since only a small number of parameters are decoupled, the increase in parameter count is minimal. Notably, experiments show that this design enables multiple tasks to promote each other.
@@ 4.3 Training and Inference Details
Training. SAM 2++ performs multi-task training on tracking tasks with different granularity (mask, box, point), initialized from SAM 2 base. We decoupled the memory-related modules, including creating a separate copy of the memory encoder and implementing dedicated LoRA parameters for memory attention for each task. In addition to our Tracking-Any-Granularity dataset, we adopt DAVIS 2017 [48], YoutubeVOS 2019 [60] and MOSE [19] for the mask task; LaSOT [23], GOT10k [31], TrackingNet [44] and COCO [41] for the box task; TapVid Kinetics [21], PointOdyssey [71], and PerceptionTest [49] for the point task. During training, we use 8-frame sequences with up to 3 targets in the first frame, The first frame and one randomly selected frame serve as conditional frames, receiving either normal prompts or interactive prompts with equal probability. For further details, please refer to the Appendix.
Losses and optimization. For mask tracking, we combine focal and dice losses for mask prediction, L1 loss for IoU prediction, and cross-entropy loss for occlusion prediction. Box tracking extends it with ciou [72] and L1 losses for box predictions. Point tracking adds L1 loss for point predictions with soft argmax. In multi-prediction scenarios, we only supervise the task prediction with the lowest combined loss while supervising all IoU predictions. For occluded frames, denoted as $1-\mathbbm{1}_{obj}$, we skip the supervision of the task results and IoU prediction, but maintain occlusion prediction supervision. In summary, the multi-task training loss can be written as:
@@ 5 Data
BURST [1] 2914 1734 (0.60) 624240 (214) $\geq$480p 6 600157 (206) × × Semi-Automatic Multi-Task
LVOS [28] 220 351 (1.60) 126280 (574) 720p 6 156432 (711) × × Manual Long-term
LVOS v2 [29] 720 823 (1.14) 296401 (412) 720p 6 407945 (567) × × Manual Large-scale, long-term
MOSE [19] 2149 443.62 (0.21) $\sim$159600 (73) 1080p 6 431725 (201) × × Semi-Automatic Complext scenarios
YoutubeVOS-19 [60] 4453 334.8 (0.08) 120532 (27) 720p 6 197272 (44) × × Manual Large-scale
VOST [51] 713 252 (0.35) 75547 (106) 1080p 5 175913 (247) × × Semi-Automatic Object transmission
LaSOT [23] 1400 1950 (1.39) 3.52M (2506) 720p 30 × 3.52M × Manual Large-scale, long-term
@@
TrackingNet [44] 30643 8400 (0.27) 14431266 (471) 360p 30 × 14431266 × Semi-Automatic Large-scale
UAV123 [4] 123 62.5 (0.51) 112578 (915) 720p 30 × 112578 × Semi-Automatic Unmanned aerial vehicles
NfS [36] 100 26.58 (0.27) 383K (3830) 720p 240 × 383K × Manual High Frame Rate
OTB-100 [58] 100 32.8 (0.33) 59040 (590) $\geq$360p 30 × 59040 × Manual Real world
TNL2K [54] 2000 691.3 (0.35) 1244340 (622) 720p 30 × 1244340 × Manual Language-based
VastTrack [46] 50610 11664 (0.23) 4.2M (83) 480p-720p 6 × 4.2M × Manual Abundant categories
@@
PointOdyssey [71] 104 120 (1.15) $\sim$216K (2035) 540p 30 × × 49B (0.471B) Automatic Real world, long-term
TAP-Vid Kinetics [21] 1189 198.17 (0.17) 297250 (250) $\geq$720p 25 × × 4725959 (3974) Semi-Automatic Abitrary point
TAP-Vid DAVIS [21] 30 - (-) 1999 (66.6) 1080p - × × 28824 (960.8) Semi-Automatic Abitrary point
TAP-Vid RGB-Stacking [21] 50 - (-) 12500 (250) 256x256 - × × 303436 (6068.7) Semi-Automatic Abitrary point
Tracking-Any-Granularity 6000 1338.7 (0.22) 2200891 (367) mostly 720p 30 2148716 (358) 2148716 2640987 (440) Semi-Automatic Any Granularity
We developed a comprehensive dataset, termed Tracking-Any-Granularity (TAG), with annotations across three granularities: segmentation masks, bounding boxes, and key points. Our dataset contains 6,000 high-resolution videos featuring diverse scenes, objects, and challenging scenarios (e.g., occlusion, motion blur, etc.). With a three-phase data engine with model-in-the-loop annotation workflows and strict multi-stage quality checks, we ensure large-scale, high-quality, and consistent annotations. For further details, please refer to the Appendix.
@@ 5.3 Tracking-Any-Granularity Dataset
STCN [12] 45.8 45.6 45.9 62.0 58.5 65.4 50.1 46.1 54.1 70.4 65.9 75.0 76.2 72.2 80.2 32.7 22.7 82.8 81.3 85.6 78.3 86.0
AOT-SwinB [67] 54.9 55.3 54.8 73.8 70.1 77.4 60.2 56.2 64.1 78.1 73.1 83.2 80.9 76.4 85.4 39.3 29.4 85.3 84.6 89.5 79.3 87.7
DeAOT-SwinB [66] 58.5 58.0 58.6 72.1 68.5 75.8 61.7 57.6 65.9 79.6 74.8 84.4 81.6 77.3 85.9 43.0 29.4 86.4 85.4 90.3 80.6 89.3
XMem [11] 52.3 51.9 52.3 64.7 61.8 67.5 56.3 52.1 60.5 74.4 70.1 78.6 75.7 71.8 79.6 37.9 25.0 85.5 84.3 88.6 80.5 88.7
DEVA [13] 56.1 55.4 56.2 72.2 68.6 75.9 66.4 62.2 70.6 77.9 73.1 82.6 82.1 78.0 86.1 40.9 27.3 86.3 85.3 89.8 80.6 89.2
Cutie-base+ [14] 55.8 55.6 55.8 71.4 68.6 74.3 66.2 62.3 70.1 79.0 75.0 83.0 83.8 80.0 87.7 44.7 32.7 86.9 86.2 90.7 81.6 89.2
@@
OSTrack [68] 74.8 84.4 72.7 69.7 78.8 69.9 68.3 77.1 66.2 83.9 88.5 83.2 57.6 74.4 58.9 33.7 40.8 31.4
SimTrack [45] 71.1 80.5 68.1 64.1 72.4 60.5 65.8 73.7 63.7 82.3 - 86.5 54.4 70.2 53.7 34.5 40.5 30.4
@@
DropTrack [57] 76.8 86.9 74.4 71.1 80.5 72.1 70.8 80.4 69.4 83.8 88.5 83.1 58.5 75.7 60.3 37.5 45.9 36.4
GRM [24] 73.1 82.3 71.4 69.1 77.4 69.1 68.5 77.3 66.5 84.0 88.7 83.3 56.8 73.2 57.7 34.6 42.2 32.3
SeqTrack [10] 77.0 85.8 76.1 69.8 79.4 71.5 68.5 78.2 67.8 83.9 88.8 83.6 57.8 75.3 60.8 35.8 44.8 35.3
ARTrack [56] 76.8 85.8 75.7 71.1 78.7 70.9 69.9 76.9 67.1 84.2 88.7 83.5 57.9 73.9 59.6 35.7 42.1 32.4
ARTrack-V2 [3] 76.3 85.5 74.3 71.8 79.5 71.9 70.2 78.0 68.3 84.9 89.3 84.5 57.1 73.4 58.7 37.0 44.5 34.8
ROMTrack [7] 75.6 85.4 73.7 71.3 80.8 72.8 69.2 77.8 68.5 84.1 89.0 83.7 58.2 75.3 59.8 37.1 45.5 36.2
HIPTrack [6] 78.2 88.5 76.6 71.4 81.0 72.5 72.0 82.2 71.0 84.5 89.1 83.8 59.8 77.1 62.1 38.6 46.3 36.8
LoRAT [40] 75.1 84.8 74.4 70.5 79.7 68.7 72.7 82.2 74.4 83.5 87.9 82.1 58.8 76.2 61.4 38.7 41.1 37.8
Ours 80.7 89.7 77.8 78.0 85.7 81.5 78.2 86.2 82.0 86.0 90.1 87.3 59.2 73.1 61.6 55.0 65.6 60.4

### v3

@@ 1 Introduction
Video tracking has been a fundamental task in computer vision for decades, aiming to estimate the state of an arbitrary target in video sequences given its initial status. Despite sharing this core objective, the tracking domain has fragmented into several independent sub-tasks based on different target granularities, including Single Object Tracking [lasot, TrackingNet, got10k] (SOT) with bounding box, Video Object Segmentation [davis17, youtube-vos, LVOS_V1] (VOS) with precise pixel-level mask, and Point Tracking [badja, zheng2023pointodyssey, tapvid] with tiny points. This fragmentation based on state granularity has led most video tracking research to focus on a specific task and propose specialized designs only for that task. While this design trend enhances tracking performance, it limits the generalization ability of tracking models across multiple tasks and results in redundancy in both model design and parameters. To unify tasks, current unified vision models typically share feature extraction backbones while employing task-specific branches [zhu2022uniperceiver], convert those tasks into a seq2seq framework [chen2021pix2seq], or share one appearance model for either propagation or association [wang2021UniTrack, yu2024unifiedtt, Unicorn, UNINEXT, OmniTracker]. However, they choose to provide different interfaces for different tasks, rather than seeking a unified visual representation of tracking targets, and ignore the point tracking task.
Unlike them, we observe that these seemingly disparate tracking paradigms fundamentally differ primarily in their state granularity, while sharing the memory matching strategy: the model encodes the previous state into memory, and matches the current features with the stored memory when a new frame is received. Based on this strategy, we decide to unify target states at three different granularities through a uniform memory representation. Recently, Segment Anything Model 2 [ravi2024sam2], a strong foundational model, has been proposed for high-quality video object segmentation given various prompts. Due to its flexible prompt mechanism and powerful mask tracking capabilities, we extend this model to track arbitrary granularity, termed as SAM 2++.
@@ 2 Related work
Segment Anything Model. SAM [kirillov2023seganysam1] is a foundational model for high-quality segmentation given various prompts, and SAM 2 [ravi2024sam2] extends it to video with streaming memory, effectively handling motion and occlusion, and they inspire many variant models In the image domain, HQ-SAM [sam_hq] enhances segmentation quality through a High-Quality Token, SAMRefiner [lin2025samrefiner] improves fine-grained details via a noise-tolerant prompt, CAT-SAM [xiao2024cat] adopts a conditional tuning approach to adapt to specialized image domains, and SAM-Adapter [chen2023samadapter] incorporates lightweight adapters for improved downstream performance. In the video domain, SAM2Long [ding2024sam2long] employs constrained tree search to reduce error accumulation, SAMURAI [yang2024samurai] uses the Kalman filter to select motion-aware memory, while DAM4SAM [videnovic2024distractor] introduces a distractor-aware memory. SAMWISE [cuttano2024samwise] and AL-Ref-SAM-2 [huang2025unleashing] add additional prompts for more referring tasks. Despite these advances, these works remain task-specific, lacking cross-domain generalization and requiring separate implementations for each application.
Unified Vision Models. Recent years have witnessed significant progress in developing unified vision models that handle multiple tasks through shared architectures and demonstrate strong generalizability and flexibility. Pix2Seq [chen2021pix2seq] reformulates vision tasks as sequence generation problems, Uni-Perceiver [zhu2022uniperceiver] establishes unified representation spaces across modalities with shared encoders and decoders. UniTrack [wang2021UniTrack] demonstrates that video tracking tasks can be solved by a single appearance model with task-specific heads, while Unicorn [Unicorn] and UNINEXT [UNINEXT] unify various tracking paradigms through common frameworks with different representations. Despite their impressive capabilities, these unified approaches predominantly focus on object-level tasks while neglecting finer-grained tasks such as point tracking. Furthermore, they do not take into account unifying video tracking tasks with various granularities through a unified visual representation.
@@ 3 Preliminaries: Segment Anything Model 2
The Segment Anything Model (SAM) [kirillov2023seganysam1] is a milestone vision foundation model for class-agnostic image segmentation. It flexibly handles various prompts (box, point, mask) by encoding them into a unified embedding and has established an iterative data engine with model-assisted labeling to address dataset limitations. SAM 2 [ravi2024sam2] extends SAM to promptable video segmentation by introducing a streaming memory that stores previous target information and predictions. It comprises four main components: (i) a hierarchical image encoder that encodes each frame ${I}_{img}$ into image embeddings $F_{img}$, (ii) a prompt encoder, (iii) a memory mechanism (memory encoder, memory bank, memory attention), and (iv) a mask decoder for prediction.
@@
Mask Decoder. The mask decoder takes prompt embedding $\mathcal{P}_{sparse}$ and $\mathcal{P}_{dense}$, memory-conditioned image embeddings $\bar{F}_{img}\in\mathbb{R}^{C/4\times H/16\times W/16}$ (which we will explain latter), and a set of learnable tokens $\mathcal{E}_{tokens}$ as inputs. The learnable tokens contain an existence token $\varepsilon_{obj}\in\mathbb{R}^{C}$ to predict whether the target exists, an IoU token $\varepsilon_{iou}\in\mathbb{R}^{C}$ to predict the result accuracy, and multiple mask tokens $\varepsilon_{mask}^{N}\in\mathbb{R}^{N\times C}$ used to obtain $N$ mask candidates. To fuse the prompt embedding, a Two-Way Transformer $\mathbf{twTrans}$ [ravi2024sam2] processes them as:
@@ 4.1 Unified Task Input and Output Processing
Input Unification via Task-Specific Prompt. Due to the input of the three tracking tasks having inconsistent granularity, we first unify inputs with task-specific prompts for different tasks. The video object segmentation task still adopts mask input ${I}_{mask}$ as its prompt in mask form, and the single object tracking task takes its box input ${I}_{box}$ as the prompt. As for the point tracking task, expect the point coordinates ${I}_{point}$, we add a dense mask ${G}_{point}$ for additional prompt as a Gaussian map centred on the point and parameterised by sigma $\sigma$ and radius $r$ as: $G_{point}=\exp\left(-\frac{\|p-p_{0}\|^{2}}{2\sigma^{2}}\right)\cdot\mathbf{1}_{\{\|p-p_{0}\|\leq r\}}$ to highlights the point in mask form, maintaining consistency with output from Unified Decoder and source for Memory Encoder, which is better than naive {0, 1} mask. More importantly, we gradually decrease the radius and sigma during training to facilitate smoother convergence and more stable learning.
Output Unification via Unified Decoder. To unify the output of various tasks, we extended the Mask Decoder of SAM 2 as Unified Decoder, which also processes memory-conditioned image embeddings, prompt embeddings, and learnable tokens. For the SOT task, the outer box of the mask output $M^{N}_{box}$ cannot be used as task output because the complexity of the mask reduces the accuracy of the box, which focuses on the center point’s position and target scale. Instead, we add a Corner-based Head [yan2021learning_stark], $\mathbf{CornHead}$, which explicitly optimizes for the accuracy and stability of the bounding box and is widely used in the SOT task, to produce box predictions. As for the PT task, rather than direct point coordinates, we obtained point predictions in terms of mask predictions by soft-argmax operation during training or argmax operation during inference. This specific design aims for the output of the point task to be consistent with the source of memory, thereby achieving a unified encoding mask output for memory information of different granularities, and also helps to optimize model training. In summary, our unified decoder can be written as:
@@ 4.2 Task-adaptive Memory
Therefore, we propose a task-adaptive memory mechanism, which relaxes the uniformity by decoupling only the memory components: each task has its own convolutional Memory Encoder, and each applies an independent LoRA [DBLP:conf/iclr/HuSWALWWC22/lora] in the transformer-based Memory Attention. This decoupled design effectively meets the diverse needs of different tasks and avoids the performance drop seen in a fully parameter-shared model, while keeping the overall structure consistent, enhancing the multi-task processing capability of the model. Since only a small number of parameters are decoupled, the increase in parameter count is minimal. Notably, experiments show that this design enables multiple tasks to promote each other.
@@ 4.3 Training and Inference Details
Training. SAM 2++ performs multi-task training on tracking tasks with different granularity (mask, box, point), initialized from SAM 2 base. We decoupled the memory-related modules, including creating a separate copy of the memory encoder and implementing dedicated LoRA parameters for memory attention for each task. In addition to our Tracking-Any-Granularity dataset, we adopt DAVIS 2017 [davis17], YoutubeVOS 2019 [youtube-vos] and MOSE [MOSE] for the mask task; LaSOT [lasot], GOT10k [got10k], TrackingNet [TrackingNet] and COCO [coco] for the box task; TapVid Kinetics [tapvid], PointOdyssey [zheng2023pointodyssey], and PerceptionTest [patraucean2023perception] for the point task. During training, we use 8-frame sequences with up to 3 targets in the first frame, The first frame and one randomly selected frame serve as conditional frames, receiving either normal prompts or interactive prompts with equal probability. For further details, please refer to the Appendix.
Losses and optimization. For mask tracking, we combine focal and dice losses for mask prediction, L1 loss for IoU prediction, and cross-entropy loss for occlusion prediction. Box tracking extends it with ciou [zheng2020diouciou] and L1 losses for box predictions. Point tracking adds L1 loss for point predictions with soft argmax. In multi-prediction scenarios, we only supervise the task prediction with the lowest combined loss while supervising all IoU predictions. For occluded frames, denoted as $1-\mathbbm{1}_{obj}$, we skip the supervision of the task results and IoU prediction, but maintain occlusion prediction supervision. In summary, the multi-task training loss can be written as:
@@ 5 Data
BURST [burst] 2914 1734 (0.60) 624240 (214) $\geq$480p 6 600157 (206) × × Semi-Automatic Multi-Task
LVOS [LVOS_V1] 220 351 (1.60) 126280 (574) 720p 6 156432 (711) × × Manual Long-term
LVOS v2 [LVOS_V2] 720 823 (1.14) 296401 (412) 720p 6 407945 (567) × × Manual Large-scale, long-term
MOSE [MOSE] 2149 443.62 (0.21) $\sim$159600 (73) 1080p 6 431725 (201) × × Semi-Automatic Complext scenarios
YoutubeVOS-19 [youtube-vos] 4453 334.8 (0.08) 120532 (27) 720p 6 197272 (44) × × Manual Large-scale
VOST [vost] 713 252 (0.35) 75547 (106) 1080p 5 175913 (247) × × Semi-Automatic Object transmission
LaSOT [lasot] 1400 1950 (1.39) 3.52M (2506) 720p 30 × 3.52M × Manual Large-scale, long-term
@@
TrackingNet [TrackingNet] 30643 8400 (0.27) 14431266 (471) 360p 30 × 14431266 × Semi-Automatic Large-scale
UAV123 [uav123] 123 62.5 (0.51) 112578 (915) 720p 30 × 112578 × Semi-Automatic Unmanned aerial vehicles
NfS [NFS] 100 26.58 (0.27) 383K (3830) 720p 240 × 383K × Manual High Frame Rate
OTB-100 [otb] 100 32.8 (0.33) 59040 (590) $\geq$360p 30 × 59040 × Manual Real world
TNL2K [wang2021tnl2k] 2000 691.3 (0.35) 1244340 (622) 720p 30 × 1244340 × Manual Language-based
VastTrack [vasttrack] 50610 11664 (0.23) 4.2M (83) 480p-720p 6 × 4.2M × Manual Abundant categories
@@
PointOdyssey [zheng2023pointodyssey] 104 120 (1.15) $\sim$216K (2035) 540p 30 × × 49B (0.471B) Automatic Real world, long-term
TAP-Vid Kinetics [tapvid] 1189 198.17 (0.17) 297250 (250) $\geq$720p 25 × × 4725959 (3974) Semi-Automatic Abitrary point
TAP-Vid DAVIS [tapvid] 30 - (-) 1999 (66.6) 1080p - × × 28824 (960.8) Semi-Automatic Abitrary point
TAP-Vid RGB-Stacking [tapvid] 50 - (-) 12500 (250) 256x256 - × × 303436 (6068.7) Semi-Automatic Abitrary point
[…]
We developed a comprehensive dataset for training our unified model, termed Tracking-Any-Granularity (TAG), with annotations across three granularities: segmentation masks, bounding boxes, and key points. Our dataset contains 6,000 high-resolution videos featuring diverse scenes, objects, and challenging scenarios (e.g., occlusion, motion blur, etc.). With a three-phase data engine with model-in-the-loop annotation workflows and strict multi-stage quality checks, we ensure large-scale, high-quality, and consistent annotations. For more details, please refer to the Appendix.
@@ 5.3 Tracking-Any-Granularity Dataset
STCN [stcn] 45.8 45.6 45.9 62.0 58.5 65.4 50.1 46.1 54.1 70.4 65.9 75.0 76.2 72.2 80.2 32.7 22.7 82.8 81.3 85.6 78.3 86.0
AOT-SwinB [aot] 54.9 55.3 54.8 73.8 70.1 77.4 60.2 56.2 64.1 78.1 73.1 83.2 80.9 76.4 85.4 39.3 29.4 85.3 84.6 89.5 79.3 87.7
DeAOT-SwinB [deaot] 58.5 58.0 58.6 72.1 68.5 75.8 61.7 57.6 65.9 79.6 74.8 84.4 81.6 77.3 85.9 43.0 29.4 86.4 85.4 90.3 80.6 89.3
XMem [xmem] 52.3 51.9 52.3 64.7 61.8 67.5 56.3 52.1 60.5 74.4 70.1 78.6 75.7 71.8 79.6 37.9 25.0 85.5 84.3 88.6 80.5 88.7
DEVA [deva] 56.1 55.4 56.2 72.2 68.6 75.9 66.4 62.2 70.6 77.9 73.1 82.6 82.1 78.0 86.1 40.9 27.3 86.3 85.3 89.8 80.6 89.2
Cutie-base+ [cutie] 55.8 55.6 55.8 71.4 68.6 74.3 66.2 62.3 70.1 79.0 75.0 83.0 83.8 80.0 87.7 44.7 32.7 86.9 86.2 90.7 81.6 89.2
@@
OSTrack [ye2022joint_OStrack] 74.8 84.4 72.7 69.7 78.8 69.9 68.3 77.1 66.2 83.9 88.5 83.2 57.6 74.4 58.9 33.7 40.8 31.4
SimTrack [pauwels2015simtrack] 71.1 80.5 68.1 64.1 72.4 60.5 65.8 73.7 63.7 82.3 - 86.5 54.4 70.2 53.7 34.5 40.5 30.4
@@
DropTrack [wu2023dropmae_droptrack] 76.8 86.9 74.4 71.1 80.5 72.1 70.8 80.4 69.4 83.8 88.5 83.1 58.5 75.7 60.3 37.5 45.9 36.4
GRM [Gao_2023_CVPR_GRM] 73.1 82.3 71.4 69.1 77.4 69.1 68.5 77.3 66.5 84.0 88.7 83.3 56.8 73.2 57.7 34.6 42.2 32.3
SeqTrack [chen2023seqtrack] 77.0 85.8 76.1 69.8 79.4 71.5 68.5 78.2 67.8 83.9 88.8 83.6 57.8 75.3 60.8 35.8 44.8 35.3
ARTrack [wei2023autoregressiveARtrack] 76.8 85.8 75.7 71.1 78.7 70.9 69.9 76.9 67.1 84.2 88.7 83.5 57.9 73.9 59.6 35.7 42.1 32.4
ARTrack-V2 [bai2024artrackv2] 76.3 85.5 74.3 71.8 79.5 71.9 70.2 78.0 68.3 84.9 89.3 84.5 57.1 73.4 58.7 37.0 44.5 34.8
ROMTrack [cai2023robust_romtrack] 75.6 85.4 73.7 71.3 80.8 72.8 69.2 77.8 68.5 84.1 89.0 83.7 58.2 75.3 59.8 37.1 45.5 36.2
HIPTrack [cai2024hiptrack] 78.2 88.5 76.6 71.4 81.0 72.5 72.0 82.2 71.0 84.5 89.1 83.8 59.8 77.1 62.1 38.6 46.3 36.8
LoRAT [lin2024tracking_lorat] 75.1 84.8 74.4 70.5 79.7 68.7 72.7 82.2 74.4 83.5 87.9 82.1 58.8 76.2 61.4 38.7 41.1 37.8
[…]